### v2

# FormulaQA: A Question Answering Dataset for Formula-Based Numerical Reasoning

Abstract
The application of formulas is a fundamental ability of humans when addressing numerical reasoning problems. However, existing numerical reasoning datasets seldom explicitly indicate the formulas employed during the reasoning steps. To bridge this gap, we propose a question answering dataset for formula-based numerical reasoning called FormulaQA, from junior high school physics examinations.  We further conduct evaluations on LLMs with size ranging from 7B to over 100B parameters utilizing zero-shot and few-shot chain-of-thoughts methods and we explored the approach of using retrieval-augmented LLMs when providing an external formula database. We also fine-tune on smaller models with size not exceeding 2B. Our empirical findings underscore the significant potential for improvement in existing models when applied to our complex, formula-driven FormulaQA.
FormulaQA: A Question Answering Dataset for
Formula-Based Numerical Reasoning
Xiao Li and Sichen Liu and Bolin Zhu and Yin Zhu and Yiwei Liu and Gong Cheng State Key Laboratory for Novel Software Technology, Nanjing University, Nanjing, China {xiaoli.nju,sichenliu,bolinzhu,yinzhu,ywliu}@smail.nju.edu.cn gcheng@nju.edu.cn
1 Introduction
Dataset Math23K-F MAWPS-F FormulaQA
# questions 23,162 2,373 5,420
# formulas (and variants) 51 (131) 18 (46) 271 (824)
# questions requiring formula (proportion) 7,750 (33.46%) 911 (38.39%) 5,420 (100%)
Avg. # reasoning steps 1.16 1.01 2.37
Table 1: Statistics of Math23-F, MAWPS-F and our FormulaQA.
Figure 1: An example taken from FormulaQA. All the numerical values (including units) are highlighted in red font. In the formulas, we marked the calculated intermediate results with purple font. All formulas are highlighted in blue.
Numerical reasoning constitutes one of the significant forms within natural language reasoning Frieder et al. (2023). The study of numerical reasoning has seen substantial progress in recent years, largely driven by the development of LLMs OpenAI (2023); Touvron et al. (2023); Li et al. (2023c) and specialized datasets Wang et al. (2017); Dua et al. (2019); Amini et al. (2019); Cobbe et al. (2021a). Current datasets for numerical reasoning typically include simple, commonsense numerical questions that do not reflect the complexity of real-world problems. These datasets do not fully address the interpretability issue in numerical reasoning, as they often rely on implicit commonsense knowledge without explicit guidance knowledge during the reasoning process. This issue becomes particularly evident when LLMs meet hallucination Frieder et al. (2023); Bang et al. (2023). Consequently, one might naturally wonder that “What knowledge could I use to guide numerical reasoning process?”. Formulas just represent the type of knowledge that is often overlooked in research but frequently utilized in real-life applications.
Take a question from the GSM8K Cobbe et al. (2021a) dataset as an example: “A robe takes 2 bolts of blue fiber and half that much white fiber. How many bolts in total does it take?”. This example only requires the use of implicit commonsense mathematical knowledge to solve without domain-specific formula. However, in FormulaQA, we require specific formulas to guide the numerical reasoning process, such as the formula used to calculate the heat absorption of an object. We annotated formulas for each question in FormulaQA. An example of FormulaQA is shown in Figure 1. The formula-based feature makes FormulaQA a more challenging dataset for developing systems that can tackle real-world numerical reasoning problems. In fields such as mathematics and physics, formulas serve as an important vessel for representing domain knowledge. However, existing datasets scarcely consider the incorporation of formulas into numerical reasoning.
Recently, Liu et al., 2023 constructed two formula-based datasets, Math23K-F and MAWPS-F. However, the formulas in these datasets primarily consist of commonsense formulas (such as total_amount = unit_amount $\times$ total_number), and only 33.5% and 38.4% of the questions in each dataset, respectively, require the use of formulas.
To address this gap, we constructed a question answering dataset for numerical reasoning that requires the use of formulas called FormulaQA. We collected questions requiring formula-based numerical reasoning from Chinese junior high school physics examinations. With the combined efforts of manual annotation and assistance from LLMs, we annotated each question with an explanation text, a final answer, and relevant formulas (including formula structures, parameter names, symbols, numerical values, and units) and built a formula database. The formula database functions as an external knowledge base, which can be used to evaluate retrieval-based/augmented systems. In Table 1, we compare FormulaQA with two existing formula-based datasets. In comparison to Math23K-F and MAWPS-F, FormulaQA contains a larger number of formulas (272), whereas the other two datasets contain 51 and 18 formulas. Additionally, all questions in FormulaQA require the use of formulas. The higher average number of reasoning steps (2.37 vs. 1.16/1.01) implies that FormulaQA is more challenging and better suited for evaluating existing models as a multi-step reasoning task.
We evaluated LLMs ranging from 7B to >100B parameters, as well as fine-tuned smaller models such as mT5-base/large Xue et al. (2021) and Qwen-1.8B Bai et al. (2023) on FormulaQA. We also attempted to train an encoder for formula retrieval and experimented with retrieval-augmented generation models. Our evaluation findings show that the best existing models only achieve an accuracy of around 70%, which is not yet satisfactory, indicating that there is still significant room for exploration in formula-based numerical reasoning.
Our contributions can be summarized as follows:
  • We construct a formula-based numerical reasoning dataset FormulaQA, with comprehensive annotations for each question.
  • We conduct evaluations on LLMs of various sizes, fine-tuned smaller models, and retrieval-augmented generative models, establishing a strong baseline for future research.
Dataset and code are available on https://github.com/nju-websoft/FormulaQA under the Apache License 2.0.
2 Related Work
2.1 Numerical Reasoning Datasets
Numerical reasoning is one of the fundamental capabilities of natural language reasoning. The study of numerical reasoning in natural language has existed for several years. Numerous datasets, such as DROP Dua et al. (2019), GSM8K Cobbe et al. (2021b), TSQA Li et al. (2021) and MATH Hendrycks et al. (2021), have introduced natural language numerical reasoning. Another line of research focusing on numerical reasoning in natural language is math word problem (MWP). MWP tasks typically provide a short passage (i.e., the question) and require the generation of an arithmetic expression that can compute the answer. Representative datasets include MAWPS Koncel-Kedziorski et al. (2016), Math23K Wang et al. (2017), MathQA Amini et al. (2019), etc.
The recently introduced datasets Liu et al. (2023) Math23K-F and MAWPS-F require formulas for only 33.5% and 38.4% of the questions, respectively, and the formulas within these datasets are all simple commonsense formulas (e.g., total_cost=unit_cost $\times$ total_number). Our FormulaQA dataset collects questions from junior high school physics examinations, with each question accompanied by domain-specific formulas. In addition, we have also annotated a formula database for FormulaQA that can serve as an external knowledge base, used to assess retrieval-augmented systems.
2.2 Numerical Reasoning Methods
The methods for solving numerical reasoning have evolved from statistical approaches Hosseini et al. (2014); Kushman et al. (2014) to those based on rules and templates Shi et al. (2015); Wang et al. (2019) and further to methods based on deep learning models Gupta et al. (2019); Chen et al. (2022); Kim et al. (2022); Li et al. (2023a). In the past two years, with the rapid development of LLMs, LLMs have demonstrated strong capabilities in resolving numerical reasoning questions. Consequently, several methods aimed at enhancing the reasoning abilities of LLMs have been proposed, including the notable Chain of Thoughts (CoTs) method Wei et al. (2022), along with many subsequent variant approaches Kojima et al. (2022); Wang et al. (2022); Zhou et al. (2022); Li et al. (2023b).
We have established representative existing methods as baselines for FormulaQA, including fine-tuning on smaller-scale models with fewer than 2 billion parameters, such as mT5-base/large, and Qwen-1.8B. Additionally, we have applied zero-shot and few-shot CoTs prompting methods to LLMs ranging from 7 billion to over 100 billion parameters. Given that existing numerical reasoning methods scarcely consider retrieval-based approaches, as the existing numerical reasoning datasets mostly rely on implicit commonsense knowledge, in FormulaQA, we have trained a specialized formula retriever for retrieving formulas and explored retrieval-enhanced numerical reasoning.
3 Dataset Construction
We collected raw questions from Chinese junior high school physics examinations from 2015 to the present, filtering out those that do not require formulas for solving. We then annotated the reasoning steps and corresponding formulas for each question. This process involved a combination of manual annotation and the assistance of LLMs to improve the efficiency of annotation. Each question included an explanation of the reasoning steps in natural language with a symbolic representation of the reasoning steps using formulas, including the values and units for all parameters within the formulas. Finally, we compiled all the formulas we obtained and merged those expressing the same meaning to create a formula database. We describe this process to construct FormulaQA in detail below.
3.1 Preprocessing
We crawled 18,433 junior high school physics examination questions in China from 2015 to the present from public sources, including only those with free-text answers and excluding multiple-choice and true/false questions. Each original question contains a question text and an explanation text that includes the reasoning steps. We eliminated questions requiring diagrams.
Subsequently, we filtered the questions by assessing the presence of numerical values within the explanation and confirming that the final answer was numerical. Utilizing a rule-based approach, we extracted the final numeric answer, including its corresponding unit, from the explanation. It was observed that for 487 instances, the final answer had not been accurately extracted from the explanations, prompting us to perform manual computations and annotations to determine the correct final answers for these questions. Following the preprocessing phase, we compiled an initial dataset comprising 6,306 questions.
3.2 Formula Normalization
Original explanation.
[…]
Explanation with normalized formulas.
Calculating the degree of temperature increase in water: [Degree of water temperature increase] = [Final temperature] - [Initial temperature] = 60 ℃ - 20 ℃ = 40 ℃. The degree of water temperature increase = 40 ℃.
Calculating the heat absorbed by water: [Heat absorbed by water] = [Mass of water] $\times$ [Specific heat capacity of water] $\times$ [Degree of water temperature increase] = 50 kg $\times$ 4.2 $\times$ $10^{3}$ J/(kg·℃) $\times$ 40 ℃ = 8400000 J. The heat absorbed by water = 8400000 J.
The thermal efficiency of the water heater can be obtained from: [Thermal efficiency of the water heater] = [Heat absorbed by water] / [Total electrical energy consumed] $\times$ 100% = 8400000 J / ($1\times 10^{7}$ J) * 100% = 84%. The thermal efficiency of the water heater = 84%.
Answer = 84%
Table 2: Original explanation and explanation with normalized formulas (highlighted in blue).
We found that the reasoning steps (i.e. the explanation) in the obtained raw dataset lacked a normalized format and were expressed quite casually. Some formulas mixed parameter names (e.g., “mass of water”) and symbols (e.g., “$m_{water}$”), while others simply provided calculations in numerical form without parameter names or symbols. In order to ensure that explanations were based on a normalized form of formulas, we then normalized the formula annotations in the explanations. An example can be found in Table 2. In this process, we need to identify the formulas used within the original explanations and to correct any formatting issues. Manually undertaking such tasks required significant effort. However, since the process is not open-ended, but rather structured, methodical and verifiable, we ensured the accuracy and consistency of the numerical reasoning by applying the extraction of formulas from the explanations. Once the formulas were identified, we could calculate each step to ensure that the results aligned with the corresponding answers. Any redundant formulas or missing parameters could be detected, so we could introduce the assistance of LLMs here.
To enhance the efficiency of the annotation process, during the LLM annotation phase, we adopted a coarse-to-fine annotation approach with the help of a LLM [^1] During dataset constructions, we accessed Qwen-max via API (https://help.aliyun.com/zh/dashscope/developer-reference/quick-start). Qwen-max is a LLM with over 100 billion parameters and a strong capability in understanding Chinese.. We first prompted the LLM in a few-shot manner to generate accurate explanations of the reasoning process. Then, we used few-shot prompts to guide the LLM in correcting minor errors within the normalized explanations, including formatting errors in formula annotations and inaccuracies in the parameters used during computations. Both prompts can be found in Appendix A.1.1. Next, we will provide a detailed description of this process.
Initially, we introduced the question along with its original explanation and the corresponding answer to guide the LLM through few-shot prompting to revise the original explanation. We observed that the ability of the LLM to revise explanations towards normalized explanations remains satisfactory. To assess the correctness of the revised explanations, we extracted formulas from these explanations and then computed the answer using the numbat tool [^2] https://numbat.dev. Numbat is designed for scientific computations with support for physical dimensions and units.. In addition to providing explanations, we also required the LLM to present the values, symbols, and units of each parameter in the formulas in the form of a table. An example is shown in Figure 1.
At this stage, we used a program to check the correctness of the formula format in the explanations, including whether there were omissions in parameter names, parameter symbols, and corresponding units and these issues are all correctable. Therefore, if our program detected that the LLM had not successfully generated an accurate normalized explanation, we used few-shot prompting to identify and correct these specific errors. More details can be found in Appendix A.1.1. We observed that the questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, unclear question formulation, and so on. Some examples of these questions can be found in Appendix A.1.2. These questions have been removed from our dataset. Following this step, our dataset contains a remaining total of 5,420 questions.
3.3 Formula Database Construction
Our next step was to construct a unified formula repository for the entire dataset. Given that parameters in the same formula can be expressed differently across various problem contexts, for instance, the two formulas “[weight of water] = [mass of water] * [gravitational acceleration]” and “[weight] = [mass] * [gravitational acceleration]” both calculate the weight of an object, we need to merge these formulas into a single representation.
We divided the construction process of the formulas database into three steps: 1) Merge the formulas through symbolic rules. 2) Merge the formulas through semantics based method. 3) Manual review and error correction. In Table 3, we present the initial number of formulas and the remaining number of formulas after each step.
Step # Formulas
Before merging 12,906
After symbolic rules based merging 1,163
After semantics based merging 439
After manual review and error correction 272
Table 3: Changes in the number of formulas after each merging step.
Symbolic rules based merging.
In this step, we merged formulas through symbolic rules. Specifically, this was achieved by comparing the structure of the formulas and the symbols contained within them. Taking this as an example of judging whether two formulas have the same structure: the formulas “$f_{1}:~{}a_{1}$=($b_{1}$+$c_{1}$)/$d_{1}$”, “$f_{2}:a_{2}$=($b_{2}$+$c_{2}$)/$d_{2}$” and “$f_{3}:b_{1}$=$a_{1}$*$d_{1}$-$c_{1}$” have the same structure because $f_{2}$ can be derived from $f_{1}$ by renaming parameters, and $f_{3}$ can be obtained from $f_{1}$ by moving $b_{1}$ to the left side of the equation. Moreover, in physics, certain physical quantities are conventionally represented by specific symbols. For example, the mass of an object is often denoted by “$m$” and the density of an object is frequently represented by the symbol “$\rho$”. Subscripts are then used to distinguish which specific object the physical quantity refers to, such as “$\rho_{water}$” for the density of water. For any two formulas, we first computed all transformations of a formula to obtain a set of all variants of the formula. Then, we compared the formula structures in the two sets to determine if two formulas were structurally equivalent. If they shared the same structure, we then compared whether the symbols, with subscripts removed, were identical. If they were, we considered these two formulas to be mergeable. When merging, we retained the parameter with the shorter length from the two.
Semantics based merging.
In the symbolic rules based merging process, the semantic information of the parameter names was neglected. This led us to perform merges grounded on the semantics of the parameter names. For instance, two formulas that were not merged during the symbol fusion stage, “[density] = [mass] / [volume]” and “[density of water ] = [mass of water] / [volume of water]”, can actually be merged. We would carry out the merging of these two formulas based on the semantic information of the parameter names (for example, "density" and "density of water" are semantically similar). For formulas with identical structures, we tokenized each pair of corresponding parameters to create two sets of words [^3] We used jieba: https://github.com/fxsjy/jieba.. When the two sets overlapped, the parameters were considered to have semantic connection, and the formulas became candidates for merging. Utilizing this approach, we identified a list of potential formula mergers and then consulted the LLM for a thorough evaluation of each pair. The prompts can be found in Appendix A.1.3.
Manual review and error correction.
Upon completing the aforementioned merging process, we manually inspected the correctness of the merge, rectified instances where errors occurred in the merging, and manually merged formulas that were overlooked by the LLM.
4 Experiments Setup
In this section, we explore several methods to address the questions within FormulaQA. These include prompting LLMs using zero-shot and few-shot chains of thoughts (CoTs, Wei et al., 2022; Kojima et al., 2022), fine-tuning smaller models with less than 2 billion parameters, and training a formula retriever to retrieve formulas which are then incorporated into the LLM prompts.
4.1 Dataset Split
We divided FormulaQA into training, id (in-distribution) test, and ood (out-of-distribution) test, comprising 4,608, 421 and 391 questions, respectively. We required that all formulas in the id test must appear in the training set, whereas in the ood test, each question involves at least one formula that has not been seen in the training set. This division is designed to evaluate the generalization capabilities of fine-tuned models on formulas they have not previously encountered.
4.2 LLMs
Following Kojima et al., 2022, we incorporated the phrase “Let’s think step by step” into the zero-shot prompt to guide LLMs in generating the reasoning steps. For the few-shot setting, we randomly sampled five questions from the train set to serve as examples for in-context learning. Each example includes the question text and the reasoning steps (i.e., the explanation). Examples of the prompts can be found in Appendix A.2.1.
We conducted experiments on GPT-3.5-turbo, GLM4, and Qwen-max, with each of these models having over 100 billion parameters. We also evaluated on Llama2-7B Touvron et al. (2023), Qwen-7B/14B Bai et al. (2023), InternLM2-7B/20B Team (2023), including the base and chat versions of these models. We followed the common practice that few-shot experiments were performed on the base versions, while zero-shot experiments were conducted on the chat versions.
4.3 Fine-tuned Small Models
To examine the performance of smaller models with fine-tuning, we fine-tuned small models with sizes less than 2B parameters. We conducted experiments on mT5-base, mT5-large Xue et al. (2021), and Qwen-1.8B Bai et al. (2023) models. Given that small models have significantly weaker numerical calculation capabilities compared to LLMs Lin et al. (2020); Testolin (2023); Lu et al. (2023), we added a setting that allows the small models to perform numerical reasoning without conducting calculations themselves, instead relying on an external calculator. Specifically, one setting involves the small models directly computing the results, while the other setting requires the models to generate the formulas for the reasoning steps along with the corresponding numerical values and units of the parameters, which are then calculated with the aid of an external calculator.
4.4 Formula Retriever
We trained a formula retriever targeting the formula database on the training set. Specifically, we encoded each question using the Chinese-BERT-wwm-base Devlin et al. (2019); Cui et al. (2021) model to obtain the CLS vector of the question. Each formula in the formula database was represented by a randomly initialized vector. During training, we calculated the cosine score between the question vector and the formula vector. The retriever was then trained with in-batch negatives and contrastive learning loss Gao et al. (2021). Subsequently, for each question in the id test, we retrieved the top five formulas with the highest scores and included them in the prompt to observe the change in the performance of the LLM when provided with relevant formulas. More details can be found in Appendix A.2.2.
4.5 Implementation Details
We accessed to GPT-3.5-turbo [^4] https://platform.openai.com/docs, GLM4555https://open.bigmodel.cn/, and Qwen-max [^6] https://help.aliyun.com/zh/dashscope/developer-reference/quick-start through API calls with the default hyper-parameters. For other LLMs, we conducted experiments on NVIDIA V100-32G GPUs for 7B models, and on NVIDIA A100-80G GPUs for 14B/20B models. These LLMs generated using nucleus sampling with top_p=0.8. Models that require fine-tuning were experimented on NVIDIA V100 GPUs with Huggingface Transformers and Pytorch 2.0. For mT5-base and mT5-large, we set a learning rate of 5e-5 and a batch size of 32, testing the model after training for 50 epochs. For Qwen-1.8B, we used a learning rate of 1e-5 and a batch size of 32, and tested the model after training for 10 epochs. For the smaller models, we also utilized nucleus sampling with top_p=0.8 for generation. In the case of the smaller models, which directly outputted formulas along with corresponding parameter values and units, if the generation output contained formatting errors, we allowed the small model to retry up to 5 times until a correctly formatted output was generated. Training mT5-base, mT5-large, and Qwen-1.8B models requires 6, 12, and 12 hours respectively.
4.6 Metric
We utilized the numbat tool to evaluate the predictions generated by the model against the gold-standard answers. A prediction is deemed correct if the relative error (prediction - gold) / gold is less than 1%. We employed accuracy, which is the proportion of questions answered correctly, as our metric.
5 Experiments Results
5.1 Results of LLMs
Model Size id test ood test Avg. zero-shot CoTs GPT-3.5-turbo unknown 26.13 25.58 25.87 GLM4 >100B 65.32 65.22 65.27 Qwen-max >100B 58.67 57.80 58.25 InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs GPT-3.5-turbo unknown 32.07 29.92 31.03 GLM4 >100B 62.47 65.98 64.16 Qwen-max >100B 58.91 63.94 61.33 InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87
Table 4: Results of LLMs on FormulaQA.
The evaluation results on LLMs are shown in Table 4. GLM4 exhibited the best performance in both zero-shot and few-shot settings, surpassing the second-ranked Qwen-max by an average of 7.02 points in both id test and ood test. GPT-3.5-turbo, in comparison to GLM4 and Qwen-max, demonstrated inferior performance, which could be attributed to the likelihood that GLM4 and Qwen-max had access to a larger corpus of Chinese data during training, thereby enhancing their capability in Chinese. Among models with size not exceeding 20B, Qwen-14B demonstrated commendable performance in both zero-shot and few-shot settings. The subpar performance of Llama2 might be due to its training data being primarily in English. We also conducted few-shot testing on the chat version of LLMs with size not exceeding 20B, and the results can be found in Appendix A.2.3.
After incorporating few-shot examples, both GPT-3.5-turbo and Qwen-max demonstrated performance improvements, ranging from 0.24 to 6.14. However, similar performance changes were not observed in GLM4, possibly due to supervised fine-tuning and alignment with human preferences, which enhanced GLM4’s understanding of instructions but may have compromised its in-context learning ability.
The evaluation results indicate that even LLMs larger than 100B parameters can only achieve a maximum score of 65.98 on FormulaQA and the performance on LLMs with size not exceeding 20B is particularly unsatisfactory, underscoring the challenging nature of FormulaQA as an unresolved dataset.
5.2 Results of Fine-tuned Small Models
Model Size id test ood test Avg. w/o calculator mT5-base 960M 1.43 5.12 3.21 mT5-large 1.7B 19.00 10.74 15.02 Qwen-1.8B 1.8B 18.29 12.28 15.40 w/ calculator mT5-base 960M 33.49 30.18 31.90 mT5-large 1.7B 33.74 32.27 33.03 Qwen-1.8B 1.8B 64.78 54.45 59.81
Table 5: Results of fine-tuned small models on FormulaQA.
Table 5 shows the evaluation results for the fine-tuned small models, with and without the calculator. Recalling Section 4.1, because small models struggle with direct numerical calculations, we provided a calculator for additional support during evaluation. Except for mT5-base, which scored poorly, mT5-large and Qwen-1.8B performed better on the id test compared to the ood test, both with and without a calculator, indicating that out-of-domain formulas indeed affected model performance. However, despite the decrease in performance on the ood test, we observed that the small model still maintains generalization ability in formula prediction. We hope that dividing our test set into id test and ood test will be helpful for evaluating the generalization ability of fine-tuned models in future works.
After using calculators, the performance of the three small models improved by scores ranging from 18.01 to 44.41. Our findings revealed that focusing on reasoning while delegating numerical computations to the calculator could enhance the performance of small models, given their limited calculating capability. This discovery provides valuable insights for future deployment of numerical reasoning systems on small models.
(a) An error case caused by wrong formulas.
(b) An error case caused by wrong calculation.
Figure 2: Error cases.
5.3 Results of LLMs with Formula Retriever
Model id test (zero-shot) id test (few-shot) GPT-3.5-turbo 26.13 32.07 GPT-3.5-turbo w/ FR 22.09 33.02 GLM4 65.32 62.47 GLM4 w/ FR 70.31 65.80 Qwen-max 58.67 58.91 Qwen-max w/ FR 56.06 58.91
Table 6: Results of LLMs with Formula Retriever on FormulaQA. “w/ FR” in the table means that the model is evaluated with top-5 retrieved formulas. Prompts can be found in Appendix A.2.4.
The results of LLMs experiments utilizing the formula retriever are shown in Table 6. We found that the impact on performance varied among different LLMs when incorporating retrieved formulas into prompts. We observed a positive enhancement on GLM4, reaching the highest performance among all models, with scores of 70.31 and 65.80 on the id test with zero-shot and id test with few-shot, respectively. This indicated that incorporating a formula retriever is indeed effective, but there is still ample room to explore how to integrate the capabilities of formula retrievers with other LLMs.
5.4 Case Study and Error Analysis
We sampled 50 error cases from the id test (few-shot setting) of GPT-3.5-turbo and manually categorized the types and proportions of errors. We divided the error types into two main categories: formula errors and calculation errors. Formula errors encompass inappropriate formulas and omitted formulas, while calculation errors primarily involve inaccuracies in numerical calculation and unit errors. We found that 38% of errors were caused by incorrect formulas, while the remaining 62% were attributable to calculation errors. We provide one example for each of the two types of errors listed in Figure 2. It could be observed that FormulaQA poses challenges to existing models in terms of formula application and numerical calculation (including unit calculation and arithmetic calculation).
6 Conclusion
In this paper, we introduced FormulaQA, a question answering dataset for formula-based numerical reasoning, which consists of 5,420 questions collected from Chinese junior high school physics exams. Unlike previous datasets where only a fraction of the questions relied on commonsense formulas or lacked formula annotations, we annotated the reasoning steps with formulas for each question with both manual and LLM-assisted efforts. Furthermore, we constructed a formula database by merging formulas with similar meanings, serving as an external knowledge base for subsequent retrieval-based/augmented approaches.  We evaluated FormulaQA across various sizes of LLMs, fine-tuned small models, and retrieval-augmented LLMs, demonstrating its challenging nature as an unresolved task. Additionally, FormulaQA introduces numerical reasoning questions based on domain-specific formulas, a dimension scarcely addressed in prior datasets. Our findings indicate substantial room for improvement of existing models on formula-based numerical reasoning, thus motivating future research efforts.
Limitations
In order to inspire further research, we have also translated the dataset into English unitizing LLMs. However, we have not yet accurately assessed the quality of the translated dataset. At present, we have not released the English version of the dataset, but we will do so later after ensuring the quality of the English dataset. Additionally, our dataset is limited to the domain of physics. Although junior high school physics is not overly complex and can be understood by most people, it is still possible to explore formula-based question answering data in other domains.
Ethical Considerations
We collected examination questions from publicly available data sources, which are openly accessible and drawn from examinations across various regions in China. Our dataset focuses on the elementary physics domain and its development may contribute to natural language reasoning, automated AI systems, and physics education. We do not foresee any potential risks for now. Annotators recruited for this task were undergraduate students majoring in science and engineering from a university in China and they are all skilled in elementary physics. All annotators were compensated with course hours and academic credits.
Appendix A Appendix
A.1 Dataset Construction
A.1.1 Prompts in Formula Normalization.
The process of formula normalization is delineated into three distinct stages: the generation of natural language explanations, the extraction of the associated parameters from the explanations, and the subsequent error correction phase. The initial two stages are illustrated in Figures 3 and 4. The third stage is further splited into three specific error categories, each addressed by a dedicated prompt: input errors, where the parameters mentioned in the explanation are absent from the question; calculation errors, which occur when the calculator reports an error during the computation process; and output errors, where the final computed answer is incorrect. We provide an example here focusing on prompts for correcting calculation errors, while prompts for the other two error types can be found in our code submission. The prompts designed to correcting calculation errors are depicted in Figure 5. The entire normalization procedure employs a 6-shot prompting, an instance of which is provided herein for illustrative purposes.
A.1.2 Examples of deleted questions
The questions which remained incorrect despite multiple attempts by the LLM were of notably poor quality, including missing important reasoning steps, wrong reference answer, and so on. Here is an example of these questions in Figure 6.
A.1.3 Semantics Based Merging for Formula Database Construction
Semantics-based merging primarily employs the LLM to comprehend formulas, ascertain if two formulas are semantically equivalent, and subsequently determine whether they can be merged into a single formula. The prompt for this procedure is illustrated in Figure 7. This approach ensures that the nuanced meanings embedded within formulas are accurately captured and evaluated for potential merging, thereby enhancing the quality of formula database.
A.2 Experiments
A.2.1 Zero-shot and Few-shot Prompts
Zero-shot and few-shot prompts are shown in Figure 8.
A.2.2 Formula Retriever
Let the number of formulas in the formula database be $N$. During training, we randomly initialized a matrix $\mathbf{F}\in\mathbb{R}^{N\times d}$, where $d$ is the hidden size and the $i$-th row in $\mathbf{F}$ represented the initial representation of the $i$-th formula in formula database. We denoted a batch of questions with a batch size of $B$ as $Q=\{q_{1},q_{2},...,q_{B}\}$. The indices of the gold-standard formulas corresponding to these $B$ questions were denoted as $L=\{l_{1},l_{2},\cdots,l_{B}\}$ (i.e. the label of $q_{i}$ is $l_{i}$, where $1\leq i\leq B$).
BERT was utilized to encode each question,
$\mathbf{h}^{i}_{cls},\mathbf{h}^{i}_{1},\cdots=\mathtt{BERT}(q_{i}),1\leq i\leq B.$ (1)
Subsequently, we took the CLS vector $\mathbf{h}^{i}_{cls}$ as the representation for the $i$-th question.
We utilized in-batch negatives and contrastive learning loss,
$\mathcal{L}=-\frac{1}{B}\sum_{1\leq i\leq B}\log\frac{\exp(\cos(\mathbf{h}^{i}_{cls},\mathbf{F}_{l_{i}}))}{\sum_{1\leq j\leq B}\exp(\cos(\mathbf{h}^{i}_{cls},\mathbf{F}_{l_{j}}))}.$ (2)
Each question might correspond to multiple correct formulas, and we ensured that the same question did not appear twice in the same batch when loading the data. Based on the implementation of Chinese-BERT-wwm-base, we tested the retrieval performance on the id test set and found that Recall@5 reached 97.69%.
A.2.3 Few-shot Experiments on the LLMs of Chat Versions
Model Size id test ood test Avg. zero-shot CoTs with LLMs of chat versions InternLM-chat 20B 5.70 4.60 5.17 Qwen-chat 14B 32.07 37.60 34.73 Qwen-chat 7B 7.36 8.70 8.01 InternLM-chat 7B 7.84 7.67 7.76 Llama2-chat 7B 0.00 0.26 0.13 few-shot CoTs with LLMs of base versions InternLM-base 20B 18.29 11.25 14.90 Qwen-base 14B 44.89 36.83 41.01 Qwen-base 7B 21.14 18.16 19.71 InternLM-base 7B 9.50 8.18 8.86 Llama2-base 7B 1.43 0.26 0.87 few-shot CoTs with LLMs of chat versions InternLM-chat 20B 11.58 10.10 10.87 Qwen-chat 14B 41.38 37.93 39.72 Qwen-chat 7B 19.21 15.02 17.19 InternLM-chat 7B 10.10 7.88 9.03 Llama2-chat 7B 1.97 1.00 1.50
Table 7: Results of different versions of the LLMs with zero-shot and few-shot on FormulaQA.
In this experiment, we could compare the performance of the same version of the model under zero-shot and few-shot settings. Results are shown in Table 7. For the chat version of the LLMs, we could observe that few-shot can effectively improve model performance, with performance improvements ranging from 1.27 to 9.18 on average across id test and ood test. Comparing the performance of the base version and chat version of the same model under few-shot settings, except for minimal changes on InternLM-chat-7B and Llama2-chat-7B, the performance of the other models showed a decrease.
A.2.4 Prompts for LLMs with Formula Retriever
We added the formulas before each question in the few-shot setting. For the examples sampled from the training set, gold-standard formulas were added before each question. For the final question from the test set in both zero-shot and few-shot prompts, we included the top 5 retrieved formulas. The prompts are shown in Figure 9.
Figure 3: Prompt of the formula normalization stage 1.
Figure 4: Prompt of the formula normalization stage 2.
Figure 5: Prompt of the formula normalization stage 3: error correction for “calculation error”.
Figure 6: An example of deleted questions.
Figure 7: Prompt for semantics-based merging.
(a) Zero-shot prompt for LLMs.
(b) Few-shot prompt for LLMs.
Figure 8: Zero-shot and few-shot prompts for LLMs.
(a) Few-shot prompt for LLMs with formula retriever.
(b) Zero-shot prompt for LLMs with formula retriever.
Figure 9: Zero-shot and few-shot prompts for LLMs with formula retriever.